# NSPDI-SNN: An efficient lightweight SNN based on nonlinear synaptic pruning and dendritic integration

### v1

@@ 4.4 Introducing the Transition Gain at Various Structural Levels
The transition gain ($\bm{A}$) is the pivotal parameters governing heterogeneity in the proposed NSPDI-SNN framework. Adjusting the value of $\bm{A}$ modulates the magnitude of non-zero weights retained after reparameterization, thereby mitigating the loss of critical information induced by aggressive sparsification. To assess the effectiveness of transition gain integration across different architectural levels, we conducted experiments on three datasets using models with transition gain introduced at the synapse level ($\bm{A}$), channel (or neuron) level ($\bm{a}$), and layer level ($a$). In all experiments, we fixed the pruning thresholds as $d_{1}=99\%$ and $d_{2}=0$, ensuring consistent sparsity control for fair comparison.
Theoretically, synapse-level transition gain ($\bm{A}$) provide fine-grained control over sparsification, enabling per-neuron regulation. However, this granularity comes at the cost of a substantial increase in the number of learnable parameters. In contrast, channel-level ($\bm{a}$) and layer-level ($a$) transition gains offer coarser control, promoting global sparsity patterns with fewer additional parameters and lower computational overhead.
As illustrated in Figure 5, all three configurations demonstrate the capacity to alleviate accuracy degradation at increasing sparsity levels, highlighting the role of transition gain in enhancing sparse training dynamics. Among them, synapse-level transition gain yield the highest accuracy improvements but incur significant computational costs due to excessive parameter growth and fine-grained regulation. Although lightweight in number of parameters, layer-level transition gain exhibit inconsistent performance, particularly under moderate-to-high sparsity levels, suggesting limited regulation capacity. In contrast, the channel-level transition gain ($\bm{a}$) consistently achieves a favorable trade-off between accuracy and complexity. It provides stable performance enhancements across all evaluated datasets and demonstrates superior robustness in the medium and high sparsity regimes. Consequently, we adopt the channel-level transition gain configuration in the final NSPDI-SNN design to ensure an optimal balance between pruning effectiveness and computational efficiency.
@@
ESL-SNN Shen et al. (2023) ResNet-19 - 92.71 67.29 $2.53\times 10^{6}$ - -0.63 2
@@ 4.5 Sparsity Induced by the Pruning Threshold $d_{2}$
Figure 6 presents the relationship between the pruning threshold $d_{2}$, final network sparsity, and classification accuracy. As $d_{2}$ increases, network sparsity rises accordingly, with minimal accuracy degradation observed at moderate sparsity levels (up to approximately 98%). In this regime, the networks maintain stable performance, with only marginal reductions in accuracy relative to their dense counterparts. However, when $d_{2}$ surpasses critical thresholds (specifically, $d_{2}>0.4$ for DVS128 Gesture, $d_{2}>0.1$ for CIFAR10-DVS, and $d_{2}>0.02$ for CIFAR10), the models experience rapid performance deterioration. This trend is consistent with prior findings in sparse SNN research, where excessive pruning compromises the integrity of information flow by eliminating essential synaptic pathways. At extreme sparsity, only a limited number of active synapses remain, constraining feature propagation and degrading recognition capabilities.
@@ 4.6 Comparison with the State-of-the-Art Methods
On the DVS128 Gesture dataset, NSPDI-SNN exhibits substantial advantages. While maintaining a high average Top-1 accuracy of 97.47% and a peak accuracy of 98.26%, the model demonstrates strong robustness under extreme pruning conditions. Even at a sparsity level of 98.86%, the accuracy degradation is limited to only -2.77%, which significantly outperforms prior methods such as SCCD (Meng et al., 2023) and DC-STE (Chen et al., 2023). For example, SCCD suffers a -19.09% performance drop at 90.13% sparsity. The model achieves exceptional compactness and computational efficiency with fewer than 7K parameters (versus DC-STE). In the more challenging CIFAR10-DVS dataset, NSPDI-SNN continues to deliver competitive performance. Although the initial accuracy (78.60%) is slightly lower than that of VGGSNN (82.4%), NSPDI-SNN achieves superior robustness at extreme sparsity (98.94%), with an accuracy loss of -5.5%. Notably, NSPDI matches the accuracy of UW-UN (Shi et al., 2024) but with just 7.05K parameters (versus multi-million in UW-UN), drastically cutting computational costs and enabling real-world deployment. For the static image dataset CIFAR10, NSPDI-SNN is pruned based on the spiking ResNet19 backbone. The model achieves an initial accuracy of 96.08%, with a maximum sparsity of 98.58% and a worst-case accuracy drop of only -5.17%. This performance exceeds that of ESL-SNN Shen et al. (2023), which incurs a -1.81% accuracy loss at 91.79% accuracy while still retaining 0.63M parameters. In contrast, NSPDI compresses the parameter count to under 0.18M, significantly reducing model size and resource demands.
@@ 5 Discussion and Conclusion
Despite its promising results, several limitations remain. The introduction of NDI introduces additional computational overhead, particularly in the case of fine-grained (e.g., synapse-level) modeling. Although we mitigated this by adopting channel-level implementations, future work should explore more efficient approximations or hardware-friendly variants of nonlinear integration. Additionally, integration with complementary learning strategies such as spike-based attention, local learning rules, or meta-plasticity mechanisms may further enhance the robustness and generalization of sparse SNNs.

### v2

@@ 4.4 Introducing the Transition Gain at Various Structural Levels
The transition gain ($\bm{A}$) is the pivotal parameter governing heterogeneity in the proposed NSPDI-SNN framework. Adjusting the value of $\bm{A}$ modulates the magnitude of non-zero weights retained after reparameterization, thereby mitigating the loss of critical information induced by aggressive sparsification. To assess the effectiveness of transition gain integration across different architectural levels, we conducted experiments on three datasets using models with transition gain introduced at the synapse level ($\bm{A}$), channel (or neuron) level ($\bm{a}$), and layer level ($a$). In all experiments, we fixed the pruning thresholds as $d_{1}=99\%$ and $d_{2}=0$, ensuring consistent sparsity control for fair comparison.
Theoretically, synapse-level transition gain ($\bm{A}$) provides fine-grained control over sparsification, enabling per-neuron regulation. However, this granularity comes at the cost of a substantial increase in the number of learnable parameters. In contrast, channel-level ($\bm{a}$) and layer-level ($a$) transition gains offer coarser control, promoting global sparsity patterns with fewer additional parameters and lower computational overhead.
As illustrated in Figure 5, all three configurations demonstrate the capacity to alleviate accuracy degradation at increasing sparsity levels, highlighting the role of transition gain in enhancing sparse training dynamics. Furthermore, all achieve improved accuracy compared to the method using STDS (red dashed line). Among them, synapse-level transition gain yields the highest accuracy improvements but incurs significant computational costs due to excessive parameter growth and fine-grained regulation. Although lightweight in number of parameters, layer-level transition gain exhibits inconsistent performance, particularly under moderate-to-high sparsity levels, suggesting limited regulation capacity. In contrast, the channel-level transition gain ($\bm{a}$) consistently achieves a favorable trade-off between accuracy and complexity. It provides stable performance enhancements across all evaluated datasets and demonstrates superior robustness in the medium and high sparsity regimes. Consequently, we adopt the channel-level transition gain configuration in the final NSPDI-SNN design to ensure an optimal balance between pruning effectiveness and computational efficiency.
@@
Figure 7: Energy ratio $E_{\text{SNN}}/E_{\text{ANN}}$ of NSPDI across different $d_{2}$ values on CIFAR-10.
@@
ESL-SNN (Shen et al., 2023) ResNet-19 - 92.71 67.29 $2.53\times 10^{6}$ - -0.63 2
@@ 4.5 Sparsity Induced by the Pruning Threshold $d_{2}$
Figure 6 presents the relationship between the pruning threshold $d_{2}$, final network sparsity, and classification accuracy. As $d_{2}$ increases, network sparsity rises accordingly, with minimal accuracy degradation observed at moderate sparsity levels (up to approximately 98%). In this regime, the networks maintain stable performance, with only marginal reductions in accuracy relative to their dense counterparts. However, when $d_{2}$ surpasses critical thresholds (specifically, $d_{2}>0.08$ for DVS128 Gesture, $d_{2}>0.03$ for CIFAR10-DVS, and $d_{2}>0.01$ for CIFAR10), the models experience rapid performance deterioration. This trend is consistent with prior findings in sparse SNN research, where excessive pruning compromises the integrity of information flow by eliminating essential synaptic pathways. At extreme sparsity, only a limited number of active synapses remain, constraining feature propagation and degrading recognition capabilities.
@@
In addition, we analyzed the energy efficiency of the NSPDI-SNN under different values of $d_{2}$ on the CIFAR10 dataset. We compared the energy ratio between the SNN and its ANN counterpart ($E_{\text{SNN}}/E_{\text{ANN}}$) at the same sparsity level and network structure. The theoretical energy cost for a multiply–accumulate (MAC) operation is estimated at 4.6 pJ, while that for an accumulate (AC) operation is 0.9 pJ (Li et al., 2025). As shown in Fig. 7, when $d_{2}$ increases from 0 to 0.0125, the energy ratio $E_{\text{SNN}}/E_{\text{ANN}}$ decreases from 0.026 to 0.005. Notably, the most significant drop occurs when $d_{2}$ increases from 0 to 0.0025, where the ratio sharply falls to 0.006. As $d_{2}$ further increases, the additional computation introduced by NSPDI becomes dominant. In this case, the energy ratio is slightly affected by further pruning and a stable ratio appears for $d_{2}\geq 0.005$.
@@ 4.6 Comparison with the State-of-the-Art Methods
On the DVS128 Gesture dataset, NSPDI-SNN exhibits substantial advantages. While maintaining a high average Top-1 accuracy of 97.47% and a peak accuracy of 98.26%, the model demonstrates strong robustness under extreme pruning conditions. Even at a sparsity level of 98.86%, the accuracy degradation is limited to only -2.77%, which significantly outperforms prior methods such as SCCD (Meng et al., 2023) and DC-STE (Chen et al., 2023). For example, SCCD suffers a -19.09% performance drop at 90.13% sparsity. The model achieves exceptional compactness and computational efficiency with fewer than 7K parameters (versus DC-STE). In the more challenging CIFAR10-DVS dataset, NSPDI-SNN continues to deliver competitive performance. Although the initial accuracy (78.60%) is slightly lower than that of VGGSNN (82.4%), NSPDI-SNN achieves superior robustness at extreme sparsity (98.94%), with an accuracy loss of -5.5%. Notably, NSPDI matches the accuracy of UW-UN (Shi et al., 2024) but with just 7.05K parameters (versus multi-million in UW-UN), drastically cutting computational costs and enabling real-world deployment. For the static image dataset CIFAR10, NSPDI-SNN is pruned based on the spiking ResNet19 backbone. The model achieves an initial accuracy of 96.08%, with a maximum sparsity of 98.58% and a worst-case accuracy drop of only -5.17%. This performance exceeds that of ESL-SNN (Shen et al., 2023), which incurs a -1.81% accuracy loss at 91.79% accuracy while still retaining 0.63M parameters. In contrast, NSPDI compresses the parameter count to under 0.18M, significantly reducing model size and resource demands.
@@ 5 Discussion and Conclusion
Despite its promising results, several limitations remain. The introduction of NDI introduces additional computational overhead, particularly in the case of fine-grained (e.g., synapse-level) modeling. We alleviated this by adopting channel-level implementations, and note that the event-driven, sparse AC operations of SNNs still provide a fundamental efficiency advantage over the dense MAC operations of ANNs (especially in fully connected layers), but this advantage is partially offset in our current architecture by the use of average pooling in convolutional layers. Future work should explore more efficient approximations or hardware-friendly variants of nonlinear integration. Replacing average pooling with max pooling can be efficiently realized through logical OR operations on spike events. This modification presents a clear path to further amplify energy efficiency. Additionally, integration with complementary learning strategies such as spike-based attention, local learning rules, or meta-plasticity mechanisms may further enhance the robustness and generalization of sparse SNNs.